OpenStudy (anonymous):

how would i graph f(x)=(x)-1

OpenStudy (anonymous):

the second x is in IxI

hero (hero):

Find a grahing calculator that allows you to input absolute value.... Otherwise split into two equations f(x) = -x -1 g(x) = x -1

OpenStudy (anonymous):

i hav one but dnt know how to use it

hero (hero):

What do you have?

OpenStudy (anonymous):

a TI-84 plus graphing calculator